# Knowledge Graph Retrieval-Augmented Generation for LLM-based Recommendation

### v1

@@ 6 Limitations
Although K-RagRec effectively retrieves knowledge sub-graphs from KGs to enhance the recommendation capabilities of LLMs, it has two main limitations. Firstly, due to the GPU resource constraints, we were only able to evaluate our framework on 7b and 8b models. Additionally, we only utilize Freebase as the external KG as it is most commonly used for recommendation tasks. Therefore, in future work, we plan to extend our method to larger models to fully assess its effectiveness and scalability. We also aim to adopt other KGs, such as YAGO, DBpedia, and Wikipedia, to better understand how different knowledge sources impact the performance of the proposed method.
@@ A.1 Implementation Details
The hyper parameters used for K-RagRec and their corresponding values are shown in Table 7. The first part provides the general training setting for K-RagRec. The second part presents the details of ($\text{GNN}^{\text{Indexing}}$, $\text{GNN}^{\text{Encoding}}$). Then we list the Lora and acceleration settings. Lastly, we provide the hyper parameters for retrieval, including candidate item number $M$, adaptive retrieval policy threshold $p$, retrieve knowledge sub-graphs numbers $K$, and re-ranking knowledge sub-graphs numbers $N$. If not specified, we run all methods three times with different random seeds and report the averaged results.
@@
Movies Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. Instruction: Given the user’s watching history, select a film that is most likely to interest the user from the options. Watching history: {”History of the World: Part I”, ”Romancing the Stone”, ”Fast Times at Ridgemont High”, ”Good Morning, Vietnam”, ”Working Girl”, ”Cocoon”, ”Splash”, ”Pretty in Pink”, ”Terms of Endearment”, ”Bull Durham”}. Options: {A: ”Whole Nine Yards”, B: ”Hearts and Minds”, C: ”League of Their Own”, D: ”Raising Arizona”, E: ”Happy Gilmore”, F: ”Brokedown Palace”, G: ”Man Who Knew Too Much”, H: ”Light of Day”, I: ”Tin Drum”, J: ”Blair Witch Project”, K: ”Red Sorghum”, L: ”Flintstones in Viva Rock Vegas”, M: ”Anna, N: Roger & Me” O: ”Land and Freedom”, P: ”In Love and War”, Q: ”Go West”, R: ”Kazaam”, S: ”Thieves”, T: ”Friends & Lovers”}. Select a movie from options A to T that the user is most likely to be interested in.
@@
Books Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. Instruction: Given the user’s reading history, select a book that is most likely to interest the user from the options. Watching history: {”Practice Makes Perfect: Spanish Verb Tenses”, ”NTC’s Dictionary of Common Mistakes in Spanish”, ”Streetwise Spanish Dictionary/Thesaurus”, ”Buscalo! (Look It Up!) : A Quick Reference Guide to Spanish Grammar and Usage”, ”La lengua que heredamos: Curso de espaol para bilinges”, ”Vox Diccionario De Sinonimos Y Antonimos”, ”Schaum’s Outline of Spanish Vocabulary”, ”Nos Comunicamos (Spanish Edition)”, ”The Oxford Spanish Business Dictionary”, ”Bilingual Dictionary of Latin American Spanish”}. Options: {”A: Folk and Fairy Tales, Childcraft (Volume 3)”, B: ”The Waves”, C: ”Dead End Kids: Gang Girls and the Boys They Know”, D: ”Opera Stars in the Sun: Intimate Glimpes of Metropolitan Opera Personalities”, E: ”Five-Minute Erotica”, F: ”Spanish Verbs: Oxford Minireference”, G: ”Motorcycle Maintenance Techbook: Servicing & Minor Repairs for All Motorcycles & Scooters”, H: ”Father and Son: A Study of Two Temperaments (Classic, 20th-Century, Penguin)”, I: ”Manga Mania Fantasy Worlds: How to Draw the Amazing Worlds of Japanese Comics”, J: ”MCSE Designing a Windows Server 2003 Active Directory & Network Infrastructure: Exam 70-297 Study Guide and DVD Training System”, K: ”The Atmospheric Boundary Layer (Cambridge Atmospheric and Space Science Series)”, L: ”St. Augustine and St. Johns County: A pictorial history”, M: ”A Will to Survive: Indigenous Essays on the Politics of Culture, Language, and Identity”, N: ”Saved: A Guide to Success With Your Shelter Dog”, O: ”Mosaic (Star Trek Voyager)”, P: ”Fantastical Tarot: 78-Card Deck”, Q: ”American Sign Language-A Look at Its History, Structure and Community”, R: ”Warrior’s Heart (Zebra Historical Romance)”, S: ”The New Money Management: A Framework for Asset Allocation”, T: ”Megabrain”}. Select a book from options A to T that the user is most likely to be interested in.
A.2 Comparison with 10 Candidate Items
A.3 Used Prompt

### v2

@@ 6 Limitations
To the best of our knowledge, this work is a pioneering study in investigating knowledge-graph RAG for LLM-based recommendations. Therefore, we realise that this work still has the following three main limitations that can be improved in the future research:
  • Firstly, due to the GPU resource constraints, we were only able to evaluate our framework on 7b and 8b models. Therefore, in future work, we plan to extend our method to larger models to fully assess its effectiveness and scalability.
  • Additionally, we only utilize Freebase as the external KG as it is most commonly used for recommendation tasks. Thus we also aim to adopt other KGs, such as YAGO, DBpedia, and Wikipedia, to better understand how different knowledge sources impact the performance of the proposed method.
  • Lastly, designing an intelligent selective retrieval policy for LLM-based recommender systems is an important and challenging task. In this work, we propose to leverage popularity to determine which items to retrieve to improve retrieval efficiency. In the future, we will investigate more flexible mechanisms (e.g., reinforcement learning) to dynamically update policies according to changes in user interest.
7 Acknowledgements
The research described in this paper has been partly supported by General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from The Hong Kong Polytechnic University (project no. P0042693, P0048625, P0051361, P0052406, and P0052986).
@@ A.1 Implementation Details
The hyper parameters used for K-RagRec and their corresponding values are shown in Table 4. The first part provides the general training setting for K-RagRec. The second part presents the details of ($\text{GNN}^{\text{Indexing}}$, $\text{GNN}^{\text{Encoding}}$). Then we list the Lora and acceleration settings. Lastly, we provide the hyper parameters for retrieval, including candidate item number $M$, popularity selective retrieval policy threshold $p$, retrieve knowledge sub-graphs numbers $K$, and re-ranking knowledge sub-graphs numbers $N$. If not specified, we run all methods three times with different random seeds and report the averaged results.
Table 3: Basic statistics of three datasets and the KG. "Items in KG" indicates the number of items that appeared in both the KG and the dataset.
A.2 Graph Neural Networks
GNNs are a critical technique in graph machine learning and are widely employed in various graph tasks. By iteratively updating node representations through aggregating information from neighboring nodes, GNNs effectively capture the underlying topology and relational structure of graphs. Formally, a typical GNN operation can be formulated as follows:
@@ A.2 Graph Neural Networks
A.3 More Related Work
The remarkable breakthroughs in LLMs have led to their widespread adoption across various fields, particularly in the recommendations. Given powerful reasoning and generalization capabilities, many studies have actively attempted to harness the power of the LLM to enhance recommender systems (Geng et al., 2022; Bao et al., 2023; Qu et al., 2024; Wei et al., 2024; Zhang et al., 2023). For example, P5 (Geng et al., 2022) proposes an LLM-based recommendation model by unifying pre-training, prompting, and prediction for various recommendation tasks, such as sequential recommendation and rating predictions. Furthermore, Tallrec (Bao et al., 2023) fine-tunes LLM (i.e., LLaMA-7B) to align with recommendation data for sequential recommendations. To further capture higher-order collaborative knowledge and enhance the model’s ability to generalize users and items, TokenRec (Qu et al., 2024) proposes a masked vector-quantized tokenizer to tokenize users and items in LLM-based recommendations. Despite their effectiveness, these models often face challenges, such as hallucinations and the lack of up-to-date knowledge. While fine-tuning can partially mitigate these issues, it is resource-costly and time-consuming due to the massive parameters of LLMs. To overcome these challenges, we propose a knowledge retrieval augmented recommendation framework that leverages external KGs to provide reliable and up-to-date knowledge instead of costly fine-tuning.
A.4 Comparison with Existing Methods
Existing LLM-based recommender systems usually require frequent fine-tuning on specific datasets to address the lack of knowledge and hallucinations, which is time-consuming and costly. To solve these challenges and avoid costly fine-tuning, our work is the first to augment the recommendation performance of LLMs by retrieving structured data (i.e., knowledge graph). Specifically, our approach differs from existing work in the following ways:
  • K-RagRec introduces an indexing GNN to efficiently retrieve structured data to enhance the recommendation capability of LLMs. Although some GraphRAG approaches also introduce GNNs to capture higher-order neighbourhood information, they only apply the last layer of the GNN representation, leading to coarse retrieval (He et al., 2024; Mavromatis and Karypis, 2024). In contrast, our approach leverages the representation of each GNN layer to retrieve nuanced knowledge of both coarse and fine-grained graph structures from KG, achieving a more comprehensive and precise retrieval.
  • In the recommendation domain that pursues inference speed, excessive retrieval time can seriously degrade the user experience resulting in user churn. Although many studies have explored selective retrieval for RAGs (Yan et al., 2024; Jiang et al., 2023), it is still an open question to determine whether an item needs to be retrieved and to reduce the retrieval time in the recommendation domain. We propose to use popularity to decide whether an item needs to be retrieved based on power law distribution, which greatly reduces the retrieval time. Table 2 shows that K-RagRec achieves inference times close to direct inference while maintaining high recommendation accuracy.
  • Typical RAG methods usually incorporate the retrieved content into the prompt as text (Wu et al., 2023b; Baek et al., 2023). However, vanilla RAG methods rely on documents and paragraphs often introduce unnecessary noise and even harmful disturbance, which can negatively impact the accuracy and reliability of recommendations. In addition, the structural relationships between entities are overlooked in typical RAG, resulting in the sub-optimal reasoning capability of LLM-based recommender systems. We propose to incorporate the retrieved knowledge subgraphs (Knowledge-GraphRAG) into the query as a graph prompt, which facilitates LLMs to better understand the retrieved knowledge subgraphs and avoids long contexts.
A.5 Comparison with 10 Candidate Items
@@
A.6 Ablation Study Setting
To assess the impact of each module in K-RagRec, we compare the framework with four ablated variants: K-RagRec (-Indexing), K-RagRec (-Popularity), K-RagRec (-Re-ranking), and K-RagRec (-Encoding). (1) K-RagRec (-Indexing) eliminates the $\text{GNN}^{\text{Indexing}}$ and stores semantic information of PLM in the knowledge vector database. For the retrieved nodes, we extract their second-order sub-graphs as the retrieved knowledge sub-graphs. (2) K-RagRec (-Popularity) does not apply the popularity selective retrieval policy and retrieves all items from the user’s historical interactions. (3) (-Re-ranking) removes the re-ranking module and inputs the knowledge sub-graphs directly. (4) K-RagRec (-Encoding) removes the $\text{GNN}^{\text{Encoding}}$ and replaces it with a trainable soft prompt. The retrieved knowledge sub-graphs will be added to the prompt as triples (e.g., {Moonraker, film writer film, Christopher Wood (writer)}).
A.7 Generalization Study
Models Methods MovieLens-20M Amazon Book
ACC R@3 R@5 ACC R@3 R@5
LLama2 K-RagRec 0.539 0.740 0.795 0.390 0.581 0.671
Lora w/K-RagRec 0.539 0.783 0.863 0.405 0.580 0.796
LLama3 K-RagRec 0.597 0.797 0.839 0.428 0.628 0.706
Lora w/K-RagRec 0.611 0.775 0.814 0.424 0.622 0.732
QWEN K-RagRec 0.507 0.769 0.861 0.418 0.612 0.687
Lora w/K-RagRec 0.545 0.814 0.897 0.441 0.623 0.706
Table 7: Quantitative comparison of hallucination on the MovieLens-1M dataset. $\Delta$ denotes the reduction in hallucinations for K-RagRec compared to Direct Inference.
Models Direct Inference K-RagRec $\Delta$
LLama-2 39.1% 2.7% 93.1%
QWEN2 4.7% 0.9% 80.9%
A.8 Study of Hallucination
In this section, we present a qualitative analysis of hallucinations in the LLama-2-7b and QWEN2 models on the MovieLens dataset. Specifically, we include a few fictional movies in the candidate items to observe the probability of the fictional movie being recommended. We compare direct recommendations and recommendations augmented with K-RagRec, and the results are shown in Table 7. We note that K-RagRec significantly reduced hallucinations by 93.1% compared to direct inference on LLama-2. In contrast to LLama-2, QWEN2 rarely recommends fictional movies. Nevertheless, K-RagRec reduced hallucinations by 80.9%, demonstrating the effectiveness of our approach in addressing hallucinations.
Table 8: Performance comparison of different KG RAG-enhanced LLM recommendation methods on the cold-start dataset and QWEN2 across three metrics. The best performances are labeled in bold. ACC and R@$k$ denote Accuracy and Recall@$k$, respectively.
Methods ACC R@3 R@5
PT w/ KG-Text 0.106 0.239 0.395
GraphToken w/ RAG 0.258 0.473 0.620
G-retriever 0.185 0.384 0.488
K-RagRec 0.406 0.705 0.834
A.9 Study of Cold Start Recommendation
The cold start problem is an important issue in most recommendation research. To comprehensively evaluate our approach, we particularly design a case study to evaluate the model’s recommendation performance under the cold-start setting. Specifically, we construct a separate cold-start dataset based on the MovieLens dataset that only contains these identified cold-start items as target items. We compare K-RagRec with three KG RAG-enhanced LLM recommendation methods, and the experiment results are shown in Table 8. The results demonstrate that our proposed K-RagRec still has satisfactory performance under the cold-start recommendation scenario, highlighting the effectiveness of our framework in all the cases.
GNN Types ACC R@3 R@5
GCN (Kipf and Welling, 2016) 0.397 0.704 0.809
GAT (Velickovic et al., 2017) 0.420 0.693 0.804
Graph Transformer (Shi et al., 2020) 0.429 0.711 0.779
GraphSAGE (Hamilton et al., 2017) 0.418 0.699 0.823
A.10 Study of Four GNN Encoders
To further understand the generality of our proposed approach, we conduct a comparative study of four variants applying different GNN encoders. Specifically, we compare GCN (Kipf and Welling, 2016), GAT (Velickovic et al., 2017), Graph Transformer (Shi et al., 2020), and GraphSAGE (Hamilton et al., 2017) as four K-RagRec variants of GNN encoder. Specifically, Graph Convolutional Network (GCN) (Kipf and Welling, 2016) first introduces convolutional operations to graph-structured data. By aggregating features from neighboring nodes, GCN facilitates the learning of rich node representations. GraphSAGE (Hamilton et al., 2017) learns an aggregation function that samples and combines features from a node’s local neighborhood in an inductive setting, enabling the effective use of new nodes. Graph Attention Network (GAT) (Velickovic et al., 2017) further incorporates attention mechanisms, allowing the model to dynamically assign varying attention to neighboring nodes, thereby enhancing the focus on the most relevant information. Inspired by the success of the transformer, the Graph Transformer (Shi et al., 2020) adapts transformer architectures to graph data, enhancing the modeling of graphs, particularly textual graphs.
Table 10: Comparison of different GNN layers on the Amazon Book dataset and LLama-2-7b across three metrics. ACC and R@k denote Accuracy and Recall@k, respectively.
GNN Layers ACC R@3 R@5
3 layers 0.496 0.653 0.736
4 layers 0.506 0.690 0.780
5 layers 0.498 0.656 0.729
A.11 Study of GNN Layer Numbers
In this subsection, we evaluate the impact of the number of GNN layers on model performance. We vary the numbers of the GNN layer numbers in the range of {3, 4, 5} and test on the Amazon Book dataset and LLama-2-7b across three metrics. As observed in Table 10, the model performance first improves and then decreases as the number of GNN layers increases, and the model achieves the best results across the three metrics when setting the number of GNN layers is set to four. Therefore, a smaller number of GNN layers may not have sufficient depth to capture the intricate relationships and dependencies in the graph, leading to sub-optimal performance. On the other hand, too many layers can result in indistinguishable node representations. Thus, selecting the optimal number of GNN layers is crucial for effective model training.
(a) Accuracy
(b) Recall@5
(c) Inference Time
Figure 4: Effect of popularity selective retrieval policy threshold $p$ on MovieLens-1M and LLama-2-7b across metrics Accuracy, Recall@5 and inference time (seconds) for K-RagRec.
(a) Amazon Book: Accuracy
(b) Amazon Book: Recall@3
(c) Amazon Book: Recall@5
(a) Amazon Book: Accuracy
(b) Amazon Book: Recall@3
(c) Amazon Book: Recall@5
A.12 Used Prompt
@@
Movies Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. Instruction: Given the user’s watching history, select a film that is most likely to interest the user from the options. Watching history: {"History of the World: Part I", "Romancing the Stone", "Fast Times at Ridgemont High", "Good Morning, Vietnam", "Working Girl", "Cocoon", "Splash", "Pretty in Pink", "Terms of Endearment", "Bull Durham"}. Options: {A: "Whole Nine Yards", B: "Hearts and Minds", C: "League of Their Own", D: "Raising Arizona", E: "Happy Gilmore", F: "Brokedown Palace", G: "Man Who Knew Too Much", H: "Light of Day", I: "Tin Drum", J: "Blair Witch Project", K: "Red Sorghum", L: "Flintstones in Viva Rock Vegas", M: "Anna, N: Roger & Me" O: "Land and Freedom", P: "In Love and War", Q: "Go West", R: "Kazaam", S: "Thieves", T: "Friends & Lovers"}. Select a movie from options A to T that the user is most likely to be interested in.
@@
Books Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. Instruction: Given the user’s reading history, select a book that is most likely to interest the user from the options. Watching history: {"Practice Makes Perfect: Spanish Verb Tenses", "NTC’s Dictionary of Common Mistakes in Spanish", "Streetwise Spanish Dictionary/Thesaurus", "Buscalo! (Look It Up!) : A Quick Reference Guide to Spanish Grammar and Usage", "La lengua que heredamos: Curso de espaol para bilinges", "Vox Diccionario De Sinonimos Y Antonimos", "Schaum’s Outline of Spanish Vocabulary", "Nos Comunicamos (Spanish Edition)", "The Oxford Spanish Business Dictionary", "Bilingual Dictionary of Latin American Spanish"}. Options: {"A: Folk and Fairy Tales, Childcraft (Volume 3)", B: "The Waves", C: "Dead End Kids: Gang Girls and the Boys They Know", D: "Opera Stars in the Sun: Intimate Glimpes of Metropolitan Opera Personalities", E: "Five-Minute Erotica", F: "Spanish Verbs: Oxford Minireference", G: "Motorcycle Maintenance Techbook: Servicing & Minor Repairs for All Motorcycles & Scooters", H: "Father and Son: A Study of Two Temperaments (Classic, 20th-Century, Penguin)", I: "Manga Mania Fantasy Worlds: How to Draw the Amazing Worlds of Japanese Comics", J: "MCSE Designing a Windows Server 2003 Active Directory & Network Infrastructure: Exam 70-297 Study Guide and DVD Training System", K: "The Atmospheric Boundary Layer (Cambridge Atmospheric and Space Science Series)", L: "St. Augustine and St. Johns County: A pictorial history", M: "A Will to Survive: Indigenous Essays on the Politics of Culture, Language, and Identity", N: "Saved: A Guide to Success With Your Shelter Dog", O: "Mosaic (Star Trek Voyager)", P: "Fantastical Tarot: 78-Card Deck", Q: "American Sign Language-A Look at Its History, Structure and Community", R: "Warrior’s Heart (Zebra Historical Romance)", S: "The New Money Management: A Framework for Asset Allocation", T: "Megabrain"}. Select a book from options A to T that the user is most likely to be interested in.